. Find the limits lim (π‘₯,𝑦)β†’(2,βˆ’4) 𝑦+4 π‘₯ 2π‘¦βˆ’π‘₯𝑦+4π‘₯ 2βˆ’4οΏ½


1
Expert's answer
2021-07-13T05:15:23-0400

We need to find "lim(x,y)\to" "(2,-4)" "\frac {(y+4x)} {(2y-xy+4x)}"


We will do direct substitution to find the limit .

"=" "\frac{(-4+4(2))} {(2(-4)-2(-4)+4(2))}"


"=\frac {(-4+8)} {(-8+8+8)}"


"=\frac 4 8" "=\frac 1 2" (ans)